Gallium has two isotopes. One isotope is 69-Ga and has a relative abundance of 60.50%. The relative atomic mass of gallium is 69.70 amu. Find the relative isotopic mass of the other isotope.
1
Expert's answer
2020-04-13T02:34:49-0400

Ar(Ga)=0.3950*x + 0.6050*69=69.70;

x=(69.70-41.745)/0.3950=70.77

Answer: 70.77 amu